OpenStudy (anonymous):

what is the diameter of this circle (x+4)^2 + (y-9)^2=18

OpenStudy (mathstudent55):

The equation of a circle with center at (h, k) and radius r is: \( (x - h)^2 + (y - k)^2 = r^2 \) What is \(r^2\) in your equation?

OpenStudy (anonymous):

18

OpenStudy (mathstudent55):

Right. Now notice that \(r^ 2 =18\), so what is r, the radius?

OpenStudy (anonymous):

4.24

OpenStudy (anonymous):

wow i feel dumb now lol

OpenStudy (mathstudent55):

Don't feel dumb. Since \(r^2 = 18\), then \(r = \sqrt{18} \). You were asked for the diameter. the diameter is twice the radius, so \(d = 2\sqrt{18} \) Which can be simplified to: \(d = 6\sqrt{2} \approx 8.49\)

OpenStudy (anonymous):

thank you